# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks for Image Analysis

### v3

@@ 1 Introduction
Although these models may achieve similar accuracy on a given test dataset, their performance can vary significantly under different operational conditions. Differential testing helps identify these variations by generating inputs that expose where one model might outperform the others, providing deeper insights that go beyond simple accuracy metrics. Similarly, in model compression, large DNN models are often compressed using techniques such as quantization, pruning, or knowledge distillation to make them suitable for deployment in resource-constrained environments, such as mobile devices or embedded systems. Differential testing is essential here to ensure that the compressed models retain the accuracy of the original uncompressed versions. By generating inputs that highlight behavioral differences between the compressed and original models, differential testing allows developers to understand better the trade-offs involved in compression and to ensure that the model’s critical performance characteristics are preserved. For regression analysis, when DNN models are updated or re-trained, differential testing identifies and generates inputs where the new version performs worse than the original or vice versa, ensuring that updates do not compromise the accuracy of the model. Last, when using ensembles of models, a voting mechanism is required, and one needs to be able to learn when to trust which model’s output.
However, the differential testing process presents several challenges. First, the input subspace where highly accurate models mispredict is often extremely limited, within a huge input space. This makes it challenging to find the specific inputs that will trigger behavioral differences, especially in models designed to perform well across a wide range of scenarios. Efficient and effective techniques are thus required to reveal meaningful such discrepancies. Second, many models are treated as black-boxes, where only the inputs and outputs are visible without access to the internal workings of the model. This lack of access complicates the generation of test inputs since the internal decision-making processes remain hidden. Third, limited testing budgets constitute another constraint, particularly in scenarios where collecting, labeling, and validating test input is time-consuming and resource-intensive. In cases where models are deployed in environments that require expensive simulators or manual validation, testing costs can quickly escalate, further limiting the resources available for comprehensive testing.
@@
Several test generation approaches for the differential testing of DNN models have been proposed in the literature [4, 5, 6, 7, 8, 9, 10], but they often come with key limitations. Many of these methods are white-box [4, 5, 6, 7], requiring access to the internal structure and parameters of the DNN models under test. This restricts their applicability to proprietary models, where such internal access is typically unavailable due to privacy or intellectual property concerns. Some other approaches [6, 7] assume structural similarities between the models under test, limiting their applicability on DNN models of different architectures. Black-box differential testing methods [8, 9, 10], on the other hand, often rely on customized mutation operators applied to seed inputs to generate triggering inputs. However, these methods are highly dependent on the availability and quality of the seed data, restricting the generation process to variations of existing inputs.
@@
To evaluate DiffGAN, we constructed a benchmark consisting of eight pairs of models (16 DNN models) trained on two widely used image classification datasets. We compared its performance to DRfuzz [10], a SOTA differential testing baseline assuming the same testing budget. Our experimental results demonstrate that DiffGAN significantly outperforms the baseline, generating on average four times more triggering inputs. Moreover, these inputs include a greater number of valid images and exhibit higher diversity, enabling a more comprehensive assessment of the models’ behaviors.
We further demonstrate an application of the triggering inputs generated by DiffGAN through the development of a machine learning-based selection mechanism. This system uses a machine learning model to dynamically select between two classifiers, predicting which one will perform better based on the input characteristics. Our results show that training the selection model with inputs generated by DiffGAN leads to significantly more accurate selections compared to those trained with inputs from the baseline approach.
@@
    We propose DiffGAN, a novel, black-box test generation approach that leverages GANs and NSGA-II to find triggering inputs for differential testing of DNN models, enabling an exhaustive analysis of behavioral disagreements between models with similar accuracy levels.
@@ 2.1 Motivations
Differential testing is crucial across a variety of practical applications such as model selection, voting systems, model compression, model regression, and detecting model reuse [8, 9, 11, 10, 12]. For example, consider the context of model selection. In practice, organizations often face the challenge of choosing the most suitable model during the training phase, prior to inference time. This decision process involves choosing from multiple DNN models, each one possibly trained by different providers for the same task, using different model architectures or training datasets, and demonstrating similar accuracy on a testing dataset. In such a scenario, the ability to generate inputs that clearly indicate differences in model behavior is invaluable to guide model selection.
@@
Differential testing is also crucial in the context of model compression, a process where large DNN models are transformed into more compact versions through techniques such as quantization, pruning, and knowledge distillation [13]. This transformation is aimed at deploying these models in resource-constrained environments while preserving as much of their original accuracy as possible. In these scenarios, differential testing is invaluable as it helps identify how the behavior of the compressed model deviates from the original. By generating triggering inputs that expose these differences, developers can better understand the trade-offs involved in the compression process.
@@
Additionally, differential testing is helpful in detecting model reuse [12]. In industries where proprietary models are crucial assets, it is essential to verify that deployed models are original. Differential testing can identify if a model presented as new is actually derived from or modified from an existing model, thus protecting intellectual property and ensuring the authenticity of models.
@@ 2.2 Challenges
Generating triggering inputs to reveal behavioral disagreements between DNN models under test introduces three main challenges. First, uncovering discrepancies in highly accurate models is inherently difficult due to the very small input subspace leading to mispredictions. Because they are designed to perform well under a wide range of conditions, we require effective and efficient techniques to identify behavioral disagreements. Second, in many real-world scenarios, we do not have access to the internal information of DNN models. Indeed, such models are often proprietary and deployed on servers, accessible only through cloud APIs. This limitation constrains us to using black-box techniques to generate triggering inputs, as we only have access to the inputs and outputs of the DNN models. Finally, the testing budget is often limited, particularly when models (1) require extensive manual labor to label and validate triggering inputs, or (2) operate on expensive simulators or physical devices. These simulators usually provide a controlled environment that replicates true operational scenarios, allowing for rigorous testing without the costs and risks associated with testing models deployed on expensive physical devices. Within a realistic testing budget, we must effectively generate triggering inputs that reveal diverse behavioral disagreements between the models under test.
@@ 2.3 Assumptions
Black-Box Models: We assume that our DNN models are black-box. The internal DNN workings, such as intermediate computations and gradients, are inaccessible. This assumption is crucial, especially in scenarios involving models received from third-party providers or models trained via federated learning, where internal model details are naturally obscured to protect privacy and intellectual property. Moreover, frameworks such as TensorFlow Lite and ONNX Inference, commonly used for model quantization, usually offer only end-to-end inference APIs. These frameworks do not provide access to intermediate model data, emphasizing efficiency in inference [14].
@@
No Access to the Training Data: We assume that we do not have access to the original training datasets of the models under test. This is a common scenario in many real-world situations where models are proprietary or when model testing is provided as a service for example. Consequently, our ability to generate inputs and interpret model behavior is based solely on observations from models’ outputs and the currently available testing data, rather than insights from the training process. Similar Accuracy Levels: We assume that each pair of models under test yields similar and high accuracy on the currently available testing dataset. Such generation warrants the effective generation of triggering inputs to reveal behavioral disagreement between the DNN models, to better assess their respective capabilities and thus enable their use in the various practical applications mentioned above.
Limited Testing Budget: We assume that our test generation approach is constrained by a limited testing budget, a valid assumption in many real-world situations when (1) DNNs are deployed on physical devices or with costly simulators in the loop, or (2) when the cost of the manual labeling or the validation of triggering inputs is high. In our study, we determine the testing budget based on the feasible execution time for test generation. We thus restrict the number of inputs to generate and analyze. Furthermore, a fixed and identical testing budget allows for a fair evaluation of alternative differential test generation approaches.
@@ 2.4 Differences from Adversarial Inputs
Adversarial samples are inputs subtly modified to mislead a specific single model by incorporating nearly imperceptible changes that may lead to incorrect predictions. Conversely, triggering inputs are designed to expose discrepancies between two versions of a model—for example the original and its compressed counterpart—by demonstrating that they produce divergent outcomes when fed with the same input. This distinction is crucial for understanding how models behave after modification or for comparing the behavior of two distinct models designed for the same task. Tian et al. [9] demonstrated that, on the MNIST dataset, only 18 out of 10,000 adversarial samples generated using various adversarial attacks served as triggering inputs for differential testing. Similarly, prior studies [15, 16] indicate that adversarial inputs can be limited in their ability to expose differences between compressed models. Given that adversarial inputs target single model weaknesses and are not suitable for revealing differences in models’ behavior, they are excluded from our differential testing analysis.
@@ 3 Approach: Reformulation as an NSGA-II Search Problem
A central problem in differential testing, particularly when evaluating DNN models that exhibit similar accuracy levels, is to generate triggering inputs that reveal behavioral disagreements between models. The generation and analysis of such inputs are crucial, as they not only reveal the specific conditions under which one model may outperform another but also enhance our understanding of the models’ operational boundaries. This insight is important for extracting the precise operational contexts in which each model is likely to fare best, thereby informing more targeted and effective application of these models in real-world scenarios. Traditional approaches, such as transformation filters and adversarial techniques, often fall short in producing inputs that introduce novel characteristics or maintain realism, thus limiting their effectiveness in the comprehensive DNN assessment[17, 18] and differential testing of DNNs. To address this challenge, we introduce DiffGAN, a differential testing approach for DNN models that leverages the capabilities of GANs and NSGA-II to drive the generation of diverse triggering inputs. The use of a GAN model for this purpose results from its proven efficacy in generating realistic and new test inputs across different DNN testing scenarios [19, 20, 21, 22]. When adequately trained, GANs can produce novel inputs that are distinct from the training data and belong to the same input domain—inputs that can be classified within the existing training labels.
In a typical GAN architecture, two deep neural networks operate in tandem: a generator, with the goal of creating new images from a random latent vector exploring the GAN’s latent space, and a discriminator, whose purpose is to determine whether a given image is produced by the generator or a direct sample from the training dataset. The ultimate goal is for the generator to produce new inputs to ensure realism and compliance with the training distribution. Note that the latent space in the context of GANs refers to a high-dimensional space that encodes a compressed representation of the data on which the GAN is trained [23, 24].
@@
We also rely on NSGA-II since it is specifically adapted to our multi-objective search problem. It is widely used in the literature and has shown good performance in solving many search-based problems in DNN testing [25, 26, 27]. Many existing works in the literature have shown the effectiveness of NSGA-II in navigating the complex latent space of GANs and solving multi-objective optimization problems [28, 29, 30]. Unlike algorithms such as MOSA, the Many Objective Sorting Algorithm [31], which approaches each fitness function in isolation, NSGA-II tries to find solutions that offer diverse trade-offs among multiple objectives. More specifically, our search is driven by two objectives: (1) maximizing the difference between the output probability vectors of the inputs to generate triggering inputs, and (2) maximizing the diversity of generated inputs by the GAN model. We illustrate in Figure 1 our test generation approach and describe in the following the different steps of our approach along with the different components of our search algorithm.
@@ 3.1 GAN Training Dataset
To effectively train our GAN model, we use the testing dataset available for the models under test. Indeed, access to the original training datasets of these models is often restricted in many real-world scenarios, particularly with proprietary models, hence our reliance on testing datasets. To enrich the diversity of the GAN’s training data, we enhance the testing dataset of the models under test with five realistic image transformations, encompassing operations such as rotation, shifting, shearing, and adjustments to the images’ brightness and blurriness. These transformations are applied to every image in the dataset, following the same procedures and parameters established by Shorten et al. [32], which serve to ensure (1) the validity of the generated inputs and (2) labels preservation by these transformations. Consequently, the augmented dataset feeding into the GAN’s training process is expanded to five times its initial size, due to the application of five distinct image transformations on each original test input. Finally, we merge the original testing dataset with the augmented data to build the final GAN training dataset. We should note that training a separate GAN (or its generator) is not needed for each pair of models under test. The GAN is trained on the shared test dataset of the models under test, augmented with data transformations. Since the test dataset remains consistent across model pairs operating within the same domain, a single GAN can be reused for all model pairs tested within that domain.
@@ 3.2 GAN Model Training
We use a GAN model [23] in our approach which encompasses a generator to produce new images and a discriminator whose role is to distinguish between the generated images and real images from the GAN’s training dataset. There are unique challenges in evaluating generators in GANs, which differ fundamentally from standard DNNs. In the realm of traditional DNNs, models’ predictions across various tasks are compared against a ground truth, which consists of the actual outcomes or labels. However, generative models, such as the generator component of a GAN, follow a different paradigm. Their output is not a label but a new, synthetic image that did not previously exist in the dataset. As a result, traditional metrics based on accuracy against a ground truth are not applicable.
@@
The KID measures the similarity between the distributions of real and generated images by computing the maximum mean discrepancy between their feature representations [33]. Unlike FID, KID is calculated using a kernel function, making it more robust to variations in sample size. KID scores range from zero to one, the former reflecting superior quality in the images generated by the model.
These metrics were chosen for their strengths in assessing both the quality and diversity of images generated by GANs. FID is sensitive to differences in image distributions, making it effective for evaluating the overall alignment between generated and real data distributions. On the other hand, KID is robust against varying sample sizes, providing accurate results even when only a few images are available [33]. Together, FID ensures that the generated images closely match the real distribution, while KID maintains robustness across different dataset sizes, making the two metrics complementary in their ability to provide a thorough evaluation of GAN-generated images [33].

### v4

@@ 1 Introduction
Although these models may achieve similar accuracy on a given test dataset, their performance can vary significantly under different operational conditions. Differential testing helps identify these variations by generating inputs that expose where one model might outperform the others, providing deeper insights that go beyond simple accuracy metrics. Similarly, in model compression, large DNN models are often compressed using techniques such as quantization, pruning, or knowledge distillation to make them suitable for deployment in resource-constrained environments, such as mobile devices or embedded systems. Differential testing is essential here to ensure that the compressed models retain the accuracy of the original uncompressed versions. By generating inputs that highlight behavioral differences between the compressed and original models, differential testing allows developers to better understand the trade-offs involved in compression and to ensure that the model’s critical performance characteristics are preserved. For regression analysis, when DNN models are updated or retrained, differential testing identifies and generates inputs where the new version performs worse than the original or vice versa, ensuring that updates do not compromise the model’s accuracy. Last, when using ensembles of models, a voting mechanism is required, and one needs to be able to learn when to trust which model’s output.
However, the differential testing process presents several challenges. First, the input subspace where highly accurate models mispredict is often extremely limited, within a huge input space. This makes it challenging to find the specific inputs that will trigger behavioral differences, especially in models designed to perform well across a wide range of scenarios. Efficient and effective techniques are thus required to reveal meaningful such discrepancies. Second, many models are treated as black boxes, where only the inputs and outputs are visible, without access to the internal workings of the model. This lack of access complicates the generation of test inputs since the internal decision-making processes remain hidden. Third, limited testing budgets pose another constraint, particularly in scenarios where collecting, labeling, and validating test inputs are time-consuming and resource-intensive. In cases where models are deployed in environments that require expensive simulators or manual validation, testing costs can quickly escalate, further limiting the resources available for comprehensive testing.
@@
Several test generation approaches for the differential testing of DNN models have been proposed in the literature [4, 5, 6, 7, 8, 9, 10]. However, these approaches often come with key limitations. Many of these methods are white-box [4, 5, 6, 7], requiring access to the internal structure and parameters of the DNN models under test. This restricts their applicability to proprietary models, where such internal access is typically unavailable due to privacy or intellectual property concerns. Some other approaches [6, 7] assume structural similarities between the models under test, which limits their applicability to DNN models of different architectures. Black-box differential testing methods [8, 9, 10], on the other hand, often rely on customized mutation operators applied to seed inputs to generate triggering inputs. However, these methods are highly dependent on the availability and quality of the seed data, restricting the generation process to variations of existing inputs.
@@
To evaluate DiffGAN, we constructed a benchmark consisting of eight pairs of models (16 DNN models) trained on two widely used image classification datasets. We compared its performance to DRfuzz [10], a SOTA differential testing baseline assuming the same testing budget. Our experimental results demonstrate that DiffGAN significantly outperforms the baseline, generating on average four times more triggering inputs. Moreover, these inputs include a greater number of valid images and exhibit higher diversity, enabling a more comprehensive assessment of the models’ behaviors. We further demonstrate an application of the triggering inputs generated by DiffGAN through the development of a machine learning-based selection mechanism. This system uses a machine learning model to dynamically select between two classifiers, predicting which one will perform better based on the input characteristics. Our results show that training the selection model with inputs generated by DiffGAN leads to significantly more accurate selections compared to those trained with inputs from the baseline approach.
@@
    We propose DiffGAN, a novel, black-box test generation approach that leverages GANs and NSGA-II to identify triggering inputs for differential testing of DNN models, enabling an exhaustive analysis of behavioral disagreements between models with similar accuracy levels.
@@
  4. 4.
    To support future research and reproducibility, we have made our source code and benchmark publicly available [^1] https://github.com/zohreh-aaa/Diff_testing_GAN.
@@ 2.1 Motivations
Differential testing is crucial across various practical applications, including model selection, voting systems, model compression, model regression, and detecting model reuse [8, 9, 11, 10, 12]. For example, consider the context of model selection. In practice, organizations often face the challenge of choosing the most suitable model during the training phase, prior to inference time. This decision process involves choosing from multiple DNN models, each one possibly trained by different providers for the same task, using different model architectures or training datasets, and demonstrating similar accuracy on a testing dataset. In such a scenario, the ability to generate inputs that clearly indicate differences in model behavior is invaluable to guide model selection.
@@
Differential testing is also crucial in the context of model compression, a process where large DNN models are transformed into more compact versions through techniques such as quantization, pruning, and knowledge distillation [13]. This transformation is aimed at deploying these models in resource-constrained environments while preserving as much of their original accuracy as possible. In these scenarios, differential testing is invaluable as it helps identify how the behavior of the compressed model deviates from the original. By generating triggering inputs that expose these differences, developers can gain a better understanding of the trade-offs involved in the compression process.
@@
Additionally, differential testing helps detect model reuse [12]. In industries where proprietary models are crucial assets, it is essential to verify that deployed models are original. Differential testing can identify if a model presented as new is actually derived from or modified from an existing model, thus protecting intellectual property and ensuring the authenticity of models.
@@ 2.2 Challenges
Generating triggering inputs to reveal behavioral disagreements between DNN models under test introduces three main challenges. First, uncovering discrepancies in highly accurate models is inherently difficult due to the very small input subspace, leading to mispredictions. Because they are designed to perform well under a wide range of conditions, we require effective and efficient techniques to identify behavioral disagreements. Second, in many real-world scenarios, we do not have access to the internal information of DNN models. Indeed, such models are often proprietary and deployed on servers, accessible only through cloud APIs. This limitation constrains us to using black-box techniques to generate triggering inputs, as we only have access to the inputs and outputs of the DNN models. Finally, the testing budget is often limited, particularly when models (1) require extensive manual labor to label and validate triggering inputs, or (2) operate on expensive simulators or physical devices. These simulators usually provide a controlled environment that replicates true operational scenarios, allowing for rigorous testing without the costs and risks associated with testing models deployed on expensive physical devices. Within a realistic testing budget, we must effectively generate triggering inputs that reveal diverse behavioral disagreements between the models under test.
@@ 2.3 Assumptions
Black-Box Models: We assume that our DNN models are black-box. The internal workings of the DNN, including intermediate computations and gradients, are inaccessible. This assumption is crucial, especially in scenarios involving models received from third-party providers or models trained via federated learning, where internal model details are naturally obscured to protect privacy and intellectual property. Moreover, frameworks such as TensorFlow Lite and ONNX Inference, which are commonly used for model quantization, typically offer only end-to-end inference APIs. These frameworks do not provide access to intermediate model data, emphasizing efficiency in inference [14].
@@
No Access to the Training Data: We assume that we do not have access to the original training datasets of the models under test. This is a common scenario in many real-world situations where models are proprietary or when model testing is provided as a service, for example. Consequently, our ability to generate inputs and interpret model behavior is based solely on observations from models’ outputs and the currently available testing data, rather than insights from the training process.
Similar Accuracy Levels: We assume that each pair of models under test yields similar and high accuracy on the currently available testing dataset. Such generation warrants the effective generation of triggering inputs to reveal behavioral disagreement between the DNN models, to better assess their respective capabilities, and thus enable their use in the various practical applications mentioned above.
Limited Testing Budget: We assume that a limited testing budget constrains our test generation approach, a valid assumption in many real-world situations when (1) DNNs are deployed on physical devices or with costly simulators in the loop, or (2) when the cost of the manual labeling or the validation of triggering inputs is high. In our study, we determine the testing budget based on the feasible execution time for test generation. We thus restrict the number of inputs to generate and analyze. Furthermore, a fixed and identical testing budget allows for a fair evaluation of alternative differential test generation approaches.
@@ 2.4 Differences from Adversarial Inputs
Adversarial samples are inputs subtly modified to mislead a specific single model by incorporating nearly imperceptible changes that may lead to incorrect predictions. Conversely, triggering inputs are designed to expose discrepancies between two versions of a model—for example, the original and its compressed counterpart—by demonstrating that they produce divergent outcomes when fed with the same input. This distinction is crucial for understanding how models behave after modification or for comparing the behavior of two distinct models designed for the same task. Tian et al. [9] demonstrated that, on the MNIST dataset, only 18 out of 10,000 adversarial samples generated using various adversarial attacks served as triggering inputs for differential testing. Similarly, prior studies [15, 16] indicate that adversarial inputs can be limited in their ability to expose differences between compressed models. Given that adversarial inputs target single model weaknesses and are not suitable for revealing differences in models’ behavior, they are excluded from our differential testing analysis.
@@ 3 Approach: Reformulation as an NSGA-II Search Problem
A central problem in differential testing, particularly when evaluating DNN models that exhibit similar accuracy levels, is to generate triggering inputs that reveal behavioral disagreements between models. The generation and analysis of such inputs are crucial, as they not only reveal the specific conditions under which one model may outperform another but also enhance our understanding of the models’ operational boundaries. This insight is essential for extracting the precise operational contexts in which each model is likely to fare best, thereby informing more targeted and effective application of these models in real-world scenarios. Traditional approaches, such as transformation filters and adversarial techniques, often fall short in producing inputs that introduce novel characteristics or maintain realism, thus limiting their effectiveness in the comprehensive DNN assessment[17, 18] and differential testing of DNNs. To address this challenge, we introduce DiffGAN, a differential testing approach for DNN models that leverages the capabilities of GANs and NSGA-II to drive the generation of diverse triggering inputs. The use of a GAN model for this purpose results from its proven efficacy in generating realistic and new test inputs across different DNN testing scenarios [19, 20, 21, 22]. When adequately trained, GANs can produce novel inputs that are distinct from the training data and belong to the same input domain—inputs that can be classified within the existing training labels.
In a typical GAN architecture, two deep neural networks operate in tandem: a generator, to create new images from a random latent vector exploring the GAN’s latent space, and a discriminator, whose purpose is to determine whether a given image is produced by the generator or a direct sample from the training dataset. The ultimate goal is for the generator to produce new inputs to ensure realism and compliance with the training distribution. Note that the latent space in the context of GANs refers to a high-dimensional space that encodes a compressed representation of the data on which the GAN is trained [23, 24].
@@
We also rely on NSGA-II since it is specifically adapted to our multi-objective search problem. It is widely used in the literature and has shown good performance in solving many search-based problems in DNN testing [25, 26, 27]. Many existing works in the literature have demonstrated the effectiveness of NSGA-II in navigating the complex latent space of GANs and solving multi-objective optimization problems [28, 29, 30]. Unlike algorithms such as MOSA, the Many Objective Sorting Algorithm [31], which approaches each fitness function in isolation, NSGA-II tries to find solutions that offer diverse trade-offs among multiple objectives. More specifically, our search is driven by two objectives: (1) maximizing the difference between the output probability vectors of the inputs to generate triggering inputs, and (2) maximizing the diversity of generated inputs by the GAN model. We illustrate in Figure 1 our test generation approach and describe in the following the different steps of our approach, along with the different components of our search algorithm.
@@ 3.1 GAN Training Dataset
To effectively train our GAN model, we use the testing dataset available for the models under test. Indeed, access to the original training datasets of these models is often restricted in many real-world scenarios, particularly with proprietary models, hence our reliance on testing datasets. To enrich the diversity of the GAN’s training data, we enhance the testing dataset of the models under test with five realistic image transformations, encompassing operations such as rotation, shifting, shearing, and adjustments to the images’ brightness and blurriness. These transformations are applied to every image in the dataset, following the same procedures and parameters established by Shorten et al. [32], which serve to ensure (1) the validity of the generated inputs and (2) label preservation by these transformations. Consequently, the augmented dataset feeding into the GAN’s training process is expanded to five times its initial size, due to the application of five distinct image transformations on each original test input. Finally, we merge the original testing dataset with the augmented data to build the final GAN training dataset. We should note that training a separate GAN (or its generator) is not needed for each pair of models under test. The GAN is trained on the shared test dataset of the models under test, augmented with data transformations. Since the test dataset remains consistent across model pairs operating within the same domain, a single GAN can be reused for all model pairs tested within that domain.
@@ 3.2 GAN Model Training
We use a GAN model [23] in our approach, which encompasses a generator to produce new images and a discriminator whose role is to distinguish between the generated images and real images from the GAN’s training dataset. There are unique challenges in evaluating generators in GANs, which differ fundamentally from standard DNNs. In the realm of traditional DNNs, models’ predictions across various tasks are compared against a ground truth, which consists of the actual outcomes or labels. However, generative models, such as the generator component of a GAN, follow a different paradigm. Their output is not a label but a new, synthetic image that did not previously exist in the dataset. As a result, traditional metrics based on accuracy against a ground truth are not applicable.
@@
The KID measures the similarity between the distributions of real and generated images by computing the maximum mean discrepancy between their feature representations [33]. Unlike FID, KID is calculated using a kernel function, making it more robust to variations in sample size. KID scores range from 0 to 1, with the former reflecting superior quality in the images generated by the model.
These metrics were selected for their ability to assess both the quality and diversity of images generated by GANs. FID is sensitive to differences in image distributions, making it effective for evaluating the overall alignment between generated and real data distributions. On the other hand, KID is robust against varying sample sizes, providing accurate results even when only a few images are available [33]. Together, FID ensures that the generated images closely match the real distribution, while KID maintains robustness across different dataset sizes, making the two metrics complementary in their ability to provide a thorough evaluation of GAN-generated images [33].